OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2.

OpenStudy (anonymous):

general form of slope-intercept is y=mx+c (where m is slope & c is y-intercept) put the values.....

OpenStudy (anonymous):

which spaces do i put the values in? can you write it all out?

OpenStudy (anonymous):

@israr000

OpenStudy (anonymous):

well m=3 & y-intercept c=-2 so putting these in y=mx+c.we get y=(3)x+(-2) y=3x-2

OpenStudy (anonymous):

thankyou so much

OpenStudy (anonymous):

n.p